### v1

# A Data-Aware Fourier Neural Operator for Modeling Spatiotemporal Electromagnetic Fields

Abstract
Neural operators have recently emerged as a powerful tool for solving partial differential equations (PDEs), including Maxwell’s equations in computational electromagnetics. Prior machine learning models often fail to capture both temporal field evolution and generalization to irregular geometries. Here, we introduce a Data-Aware Fourier Neural Operator (DA-FNO) as a surrogate solver. Applied autoregressively, it predicts the temporal evolution of all field components while monitoring energy dynamics, terminating automatically once energy converges. The model generalizes to complex geometries and the optical C-band without retraining, achieving a $7.5\times$ speedup with nearly $92.5\%$ accuracy. This approach offers a potentially efficient and accurate alternative to conventional iterative solvers for C-band photonic simulations.
1 Introduction
Solving Maxwell’s equations remains a fundamental challenge in photonics. Current engineering practices predominantly rely on numerical methods such as the Finite Element Method (FEM) [1, 2], Finite Difference Frequency Domain (FDFD) [3, 4], and Finite Difference Time Domain (FDTD) [5, 6] algorithms. Although full-wave Maxwell solvers provide accuracy, their substantial computational costs restrict the scalability of applicable scenarios. Among these methods, the FDTD algorithm stands out for its explicit temporal evolution of electromagnetic fields, enabling efficient time-resolved simulations such as ultrafast optics and broadband excitation. However, its strict adherence to the Courant-Friedrichs-Lewy (CFL) stability condition imposes stringent spatiotemporal resolution constraints, leading to excessive computational costs for high-precision or large-scale simulations.
In recent years, machine learning (ML) approaches have been extensively explored for solving electromagnetic field problems, often demonstrating significantly faster computational speeds than traditional methods while maintaining acceptable accuracy. Convolutional neural networks (CNNs) with their local receptive fields and various recurrent networks, such as Recurrent CNN (RCNN) [7, 8, 9] and Long Short-Term Memory - FDTD (LSTM-FDTD) [10, 11], have been widely applied to improve the FDTD algorithm. Recent extensions also incorporate transformer architectures [12, 13, 14] to model long-range electromagnetic interactions through self-attention mechanisms. However, these conventional network architectures struggle to generalize across irregular and arbitrary geometries despite their iterative temporal prediction capabilities. In contrast to these data-driven approaches, physics-informed neural networks (PINNs) [15, 16] have emerged as an unsupervised approach that embeds physical laws into the loss function, enabling PDEs solving without labeled data. Concurrently, neural operator methods have gained attention for learning mappings between function spaces, effectively modeling the underlying PDE operators. Representative models such as Deep operator networks (DeepONets) [17] and Fourier neural operators (FNOs) [18, 19, 20, 21] have been applied to solve Maxwell’s equations, demonstrating promising results in computational electromagnetics. Nevertheless, most operator-based frameworks and PINNs focus exclusively on frequency-domain field predictions, neglecting time-domain dynamics, while the few capable of temporal outputs [18, 19] are confined to parametric variations of regular geometries.
In this work, we propose a Data-Aware Fourier Neural Operator (DA-FNO) that is applied autoregressively as a surrogate solver for 2D electromagnetic field simulations. The proposed model is capable of generalizing across unseen geometries, targeting Transverse Magnetic (TM)-polarized plane wave scattering in a simulation domain. Similar to the FDTD method, the model iteratively outputs all field components $(E_{z},H_{x},H_{y})$ over time, allowing for dynamic tracking of the scattering process. Therefore, the DA-FNO model simultaneously supports the temporal evolution of electromagnetic fields and generalization to complex and randomized geometries. In addition, it demonstrates strong wavelength generalization across the optical communication C-band. The model monitors the instantaneous energy within the domain, progressively reducing energy as the outgoing scattered waves evolve, and automatically terminates the simulation once the energy falls below a convergence criterion, $\delta$, of its historical maximum. Moreover, it eliminates CFL constraints, enabling larger time steps ($\Delta t$), achieving a $7.5\times$ speedup over the conventional FDTD method while maintaining $92.5\%$ solution accuracy when $\delta=10^{-2}$. Fig. 1 illustrates the schematic diagram of the proposed model. The FDTD simulation is performed with a time step of $\Delta t$. The DA-FNO model outputs the three field components ($E_{z},H_{x},H_{y}$) of the TM-polarized wave at every $m\Delta t$ interval, enabling faster simulation compared to the FDTD. This process continues until the energy reaches convergence.
Figure 1: Schematic diagram of the proposed model. The time step of FDTD is $\Delta t$ while that of the DA-FNO model is $m\Delta t$, enabling faster simulation.
2 Model Construction
Prior to introducing our proposed network architecture, we first provide a concise overview of the vanilla Fourier Neural Operator (the original FNO without modifications) [22]. The vanilla FNO architecture is designed to learn mappings between infinite-dimensional function spaces for solving PDEs. It is inspired by kernel integral operators for solving PDEs, analogous to Green’s function-based convolutions. For spatial problems, by leveraging Fourier transforms, the vanilla FNO translates spatial convolutions into spectral multiplications, effectively parameterizing the integral kernel through learnable linear transformations in the spectral domain. Fig. 2(a) illustrates the full architecture of the vanilla FNO, while Fig. 2(b) shows the structure of the Fourier layer used in (a). In detail, the input function $a\left(x\right)\in R^{d_{in}}$ is first lifted to a high-dimensional latent space ${v_{0}\left(x\right)\in R}^{w}$ via a shallow fully connected network $P$. Each subsequent Fourier layer processes its corresponding input function through: (1) a Fourier transform projecting the input function to spectral space, (2) a linear transformation via learnable spectral weight $R_{m}$, applied within a truncated spectral band retaining only low-frequency modes, (3) an inverse Fourier transform reconstructing spatial function, (4) a residual addition combining the layer input function and the output function from step (3), and (5) a nonlinear activation function (ReLU) applied to the residual-summed function. The output function of the Fourier layer is given by
$v_{m+1}\left(x\right)=\sigma\left(\mathcal{F}^{-1}\left(R_{m}\odot\mathcal{F}\left(v_{m}\left(x\right)\right)\right)+W_{m}v_{m}\left(x\right)\right),$ (1)
where m denotes the index of the Fourier layer and $W_{m}$ is a learnable linear transformation for the $m$-th layer. The symbol $\odot$ denotes the Hadamard product, representing element-wise multiplication. After Fourier layers, $v_{n}\left(x\right)$ is projected to the target function $u(x)\in R^{d_{out}}$. Benefiting from spectral-domain transformations, the vanilla FNO can effectively capture global spatial dependencies, making it particularly advantageous for solving PDEs compared to conventional convolution-based methods.
Figure 2: (a) The full architecture of the vanilla FNO. $a(x)$ is lifted to a higher-dimensional space by network $P$, processed through $n$ Fourier layers and projected to the target function $u(x)$ by network $Q$. (b) The architecture of the Fourier layer in (a). The output $v_{0}(x)$ from network $P$ is Fourier-transformed into the spectral space, where a spectral weight $R$ is applied along with rectangular truncation. After an inverse Fourier transform, the result is added to $v_{0}(x)$ and passed through an activation function.
Our model utilizes a Data-Aware Fourier Neural Operator (DA-FNO) within an auto-regressive framework to iteratively predict time-evolving electromagnetic fields. The main difference between the DA-FNO and the vanilla FNO is the design of the Fourier layer, which will be discussed in detail later.
A pictorial representation of our model architecture is presented in Fig. 3(a). The input tensor of the model stacks five consecutive temporal states $\left\{s_{t-4},\ldots,s_{t}\right\}$ along the channel dimension, with each state comprising electromagnetic field components $\left(E_{z},H_{x},H_{y}\right)$ for the TM-polarized wave. Spatial coordinates $\left(x,y\right)$ and the permittivity distribution $\epsilon$ are first appended to the channels to encode geometric priors. Next, the augmented input is lifted into a high-dimensional latent space via a fully connected layer $P$, and subsequently transformed through four Data-Aware Fourier layers. Then, a linear layer $Q$ maps the refined latent features to the output temporal state $s_{t+1}$, representing the electromagnetic field updates. Finally, the oldest state in the input tensor $s_{t-4}$ is discarded, and the predicted state $s_{t+1}$ is appended to form the updated input tensor $\left\{s_{t-3},\ldots,s_{t+1}\right\}$ for the next iteration. This auto-regressive process continues, with the electromagnetic energy of each newly predicted state being computed as:
$U_{t+1}=\int_{S}{\frac{1}{2}\left(\mathbf{E}\cdot\mathbf{D}+\mathbf{H}\cdot\mathbf{B}\right)dS},$ (2)
where $S$ is the simulation domain. The simulation autonomously terminates when $U_{t+1}<\delta\cdot U_{max}$, and $U_{max}$ denotes the maximum energy recorded throughout the simulation history.
Figure 3: (a) The full architecture of the DA-FNO model. The five input temporal field states $\left\{s_{t-4},\ldots,s_{t}\right\}$ is augmented by $S$ with permittivity distribution $\epsilon$ and spatial coordinates $(x,y)$, lifted to a higher-dimensional space by network $P$, processed through four DA Fourier layers and mapped to the next state $s_{t+1}$ by network $Q$. The new state and the last four temporal states of the input form a new input $\left\{s_{t-3},\ldots,s_{t+1}\right\}$, enabling the iteration to continue until energy convergence. (b) The architecture of the DA Fourier layer. The convolutional processing establishes correlations among the three field components. Data-Aware mode selection is performed in the spectral space.
Compared to the vanilla FNO model, the DA-FNO model processes input data with an additional dimensionality, which is used to represent the three field components $\left(E_{z},H_{x},H_{y}\right)$. These components are inherently coupled through the underlying physical laws according to Maxwell’s curl equations without sources:
$\nabla\times\mathbf{H}$ $=\frac{\partial\mathbf{D}}{\partial t}$ (3)
$\nabla\times\mathbf{E}$ $=-\frac{\partial\mathbf{B}}{\partial t}.$
However, when directly fed into the vanilla FNO, these physically coupled field components are treated as decoupled channels during spectral processing. As a result, the vanilla FNO fails to capture the intrinsic physical correlations and constraints dictated by the curl equations, which link the evolution of one component to the spatial variations of the others. To address this architectural limitation and enforce physics-compliant interactions, we introduce three learnable $3\times 3$ convolution kernels before the Fourier transform $\mathcal{F}$ in each Fourier layer, as shown in Fig. 3(b). Here, $T$ denotes a tensor shape transformation that maps the dimension of the field component to the dimension of the feature, while $T^{-1}$ represents its inverse. Each convolution kernel takes all three field components as input channels and produces a single output component. This design allows each field component to dynamically incorporate information from the other two through localized spatial mixing. We adopt convolution rather than other linear transformations because it effectively captures local spatial dependencies, closely resembling the computations of curl operators. Combined with the Fourier transform, the DA-FNO enables complementary learning of both local and global spatial interactions.
Figure 4: (a) Ez field distributions at $t=5\Delta t,13\Delta t$ and $21\Delta t$. (b) The mode selection regions of the vanilla FNO and DA-FNO. Average of the spectral images of all input data, normalized and displayed in logarithmic scale.
Furthermore, we modify the spectral mode selection strategy within the Fourier layers of the vanilla FNO. The vanilla FNO architecture adopts a rectangular truncation window, $\left[-k_{max},k_{max}\right]^{2}$, to retain low-frequency spectral modes. While this promotes smooth outputs and suppress high-frequency noise, it may be suboptimal for electromagnetic simulations, where field interactions with complex geometries give rise to increasingly rich high-frequency features over time. As illustrated in Fig. 4(a), the $E_{z}$ field distributions at $t=5\Delta t$$13\Delta t$, and $21\Delta t$ exhibit progressively intricate spatial features, accompanied by the emergence of high-frequency components. To address this spectral adaptability gap, we propose a data-driven spectral mode selection scheme, consisting of three steps: (1) Perform 2D Fourier transforms of spatiotemporal field slices $(E_{z},H_{x},H_{y})$ across the training dataset. (2) Apply min-max normalization to each spectrum to mitigate amplitude decay effects arising from temporal evolution. (3) Average the normalized spectra and select spectral modes in descending order of magnitude until the cumulative sum exceeds a predefined threshold $\theta$ of the total integrated magnitude. The indices of the selected modes are recorded and used to define the data-driven sampling region in each Fourier layer. Fig. 4(b) illustrates the selection regions (white areas) of the vanilla FNO and the DA-FNO at $\theta=90\%$, each selecting a comparable number of spectral modes. The final panel depicts the averaged Fourier spectrum over 2700 samples. Compared to the vanilla FNO, the selection region of the DA-FNO exhibits a better match to the spectral distribution, capturing more high-amplitude spectral modes.
Both the introduction of the convolutional processing and the refinement of the mode selection strategy are motivated by the characteristics of the input data. Specifically, the data exhibits interdependent field components, which prompts the use of convolution to capture their interactions. The intensification of high-frequency components over time necessitates a revised mode selection strategy that retains more high-frequency modes. These design choices, motivated by the underlying properties of the data, collectively define our improved FNO, which we refer to as the Data-Aware FNO (DA-FNO). In addition, we use the Scaled Exponential Linear Unit (SELU) [23] activation function instead of ReLU in the Fourier layer, which as observed in our experiments, helps reduce the prediction errors.
The detailed implementation of the network architecture, developed in PyTorch, is available in the full Python code in Ref. [24].
3 Numerical Experiments
3.1 Data Generation
We generate randomized geometries following the methodology of Ref. [20]. A $256\times 256$ matrix is first initialized with uniformly distributed random values between 0 and 1. A 2D Gaussian filter $\left(\sigma=30\right)$ is then applied to the central $250\times 250$ region of the matrix, while the peripheral regions remain zero-padded to confine structural features within the domain. The smoothed output is binarized using a threshold of 0.5: values above the threshold are set to 1 (solid material), and those below to 0 (void). To mitigate abrupt material transitions at interfaces, the binary map is down-sampled to a $128\times 128$ matrix via $2\times 2$ local averaging. This averaging operation smooths interfacial discontinuities while preserving topological randomness, enabling more physically realistic representation of arbitrary geometries. We subsequently map the $128\times 128$ matrix to a permittivity distribution using a simple mapping:
$\epsilon\left(r\right)=\left[n_{void}+\left(n_{solid}-n_{void}\right)\chi\left(r\right)\right]^{2},$ (4)
where $\chi\left(r\right)\in\left\{0,1\right\}$ denotes the randomized geometry, $n_{void}=1$ and $n_{solid}=1.4$ represent air and material respectively.
We employ an in-house FDTD solver to perform the simulations and set the convergence criterion, $\delta=10^{-4}$. The discrete permittivity distribution $\epsilon\left(r\right)$ is applied within a total-field/scattered-field (TFSF) region, which is enclosed by a perfectly matched layer (PML) to suppress spurious reflections at the domain boundaries. A TM-polarized Gaussian pulsed plane wave $\left(\lambda_{0}=1550\text{nm},\tau_{FWHM}=30fs\right)$ is launched along the principal axis to simulate the interaction with the randomized scattering geometry, as shown in Fig. 5. The simulation continues until the domain energy converges $\left(U_{t}<\delta\cdot U_{max}\right)$. The spatial and temporal resolutions are set to $\Delta x=\Delta y=40\ nm/px$ and $\Delta t=\frac{1}{c}[(\frac{1}{\Delta x})^{2}+(\frac{1}{\Delta y})^{2}]^{-\frac{1}{2}}$ to satisfy the CFL stability condition.
A total of 3,000 simulation samples are generated at a $1550\text{nm}$ central wavelength, of which 2,700 are used for training and the remaining 300 for testing. To evaluate wavelength generalization, we further use the same geometries to generate 8 additional test sets (300 samples each) for central wavelengths ranging from $1530\text{nm}$ to $1570\text{nm}$ at $5\text{nm}$ intervals, excluding the $1550\text{nm}$ training wavelength. This wavelength range lies within the optical communication C-band ($1530\text{-}1565\text{nm}$). The entire data generation process took approximately an hour. The first 200 samples are available in Ref. [24], and additional data are available upon request via email.
Figure 5: Illustration of the data generation process.
3.2 Training
The proposed DA-FNO model eliminates the CFL condition, enabling larger time-step iterations while simultaneously generating all components of the electromagnetic field at each iteration. Before training, the model’s temporal resolution is set to 15 times the FDTD time step ($\Delta t_{model}=15\Delta t_{FDTD}$, corresponding to $m=15$ in Fig. 1). We sample the full-field data $\left(E_{z},H_{x},H_{y}\right)$ from datasets at every 15th FDTD time step and scale the magnetic field components by the intrinsic impedance of free space, $Z_{0}=\sqrt{\mu_{0}/\epsilon_{0}}=120\pi\ \Omega$, to match the magnitude of the electric field. Subsequently, each data sample is normalized to the range $\left[-1,1\right]$, motivating the use of a tanhtanh activation function in the model’s output layer. These preprocessing steps promote numerical stability, balanced feature scaling, and improved training generalization.
Due to the energy convergence termination criterion ($\delta=10^{-4}$) in the model simulation, the simulation duration—and consequently the temporal length of the generated data—varies across different geometries. To facilitate batch processing and accelerate training, each sample within a mini-batch is zero-padded to a uniform length $T$, and a mask is defined to record the padded entries for subsequent loss computation. For each mini-batch, the initial five time slices serve as the input $\mathbf{x}$, while the remaining T5T-5 time slices constitute the label data $\mathbf{y}$. The model iteratively predicts the field components for T5T-5 steps, and all predicted components are concatenated to form $\mathbf{\hat{y}}$. The loss function for backpropagation is then computed using $\mathbf{y}$, $\mathbf{\hat{y}}$, and the $\mathbf{mask}$:
$\mathcal{J}(\mathbf{\hat{y},\mathbf{y},\mathbf{mask}})=\sum_{s,t}(\mathbf{mask}\odot(\frac{1}{3}\sum_{c=1}^{3}\mathcal{L}_{2}(\mathbf{\hat{y}},\mathbf{y}))),$ (5)
where $s$,$t$ and $c$ denotes the sample index, time index and component index, respectively. The $\mathbf{mask}$ is a 2D tensor of 0s and 1s, where 1 indicates valid data and 0 denotes padding, excluding the influence of padded regions during backpropagation. $\mathcal{L}_{2}(\mathbf{\hat{y}},\mathbf{y})$ represents the relative $L_{2}$ loss:
[…]
The loss function $\mathcal{J}$ computes the relative $L_{2}$ error across all field components and time steps, thereby preventing the model from discounting later time steps with lower magnitudes and encouraging more consistent prediction accuracy throughout the entire temporal sequence.
All training was conducted on NVIDIA A6000 and A100 GPUs.
3.3 Results
To evaluate the performance on training and testing datasets, we use the relative $L_{1}$ error:
$\mathcal{L}_{1}\left({\hat{y}},{y}\right)=\frac{{\left\|{\hat{y}}-{y}\right\|_{1}}}{{\left\|{y}\right\|_{1}}}.$ (7)
throughout.
[…]
We first evaluate the training performance of the DA-FNO models under different threshold values $\theta$, and compare them with that of the Conv-FNO and vanilla FNO models. The Conv-FNO model refers to the vanilla FNO with the convolutional processing module from DA-FNO, while retaining the original rectangular mode selection strategy. Notably, the number of selected modes in Conv-FNO and vanilla FNO is slightly higher than that of DA-FNO with $\theta=90\%$.
Through several training runs, we observe that initial conditions have a significant impact on the training dynamics and final performance of the models. Given this sensitivity, we conduct repeated training trials with random weight initializations to ensure that the evaluation results are representative and not biased by outlier cases. For each model, these trials are performed using the first 500 samples of the full dataset, which are partitioned into a training subset (samples 1-450) and a test subset (samples 451-500). Among all trials, three runs with stable and consistent loss descent curves are selected for comparison. All models are trained for 1000 epochs under identical hyperparameter settings.
The results are presented in Fig. 6. For DA-FNO models, increasing the threshold $\theta$ leads to a consistent downward trend in both training and test errors, indicating that retaining more spectral modes improves the model’s ability to simulate electromagnetic fields. However, when $\theta$ exceeds 0.9, the remaining modes (as shown in the spectral distribution of Fig. 4(b)) exhibit very low amplitudes, and even a slight increase in $\theta$ leads to a substantial rise in the number of selected modes. Therefore, we do not consider higher values of $\theta$.
Compared with the Conv-FNO model, the DA-FNO ($\theta=90\%$) model achieves a slightly lower minimum error in the Fig. 6, suggesting that the data-aware mode selection strategy is more effective than the rectangular strategy in capturing long-term electromagnetic dynamics. Additionally, both DA-FNO and Conv-FNO models significantly outperform the vanilla FNO models, especially in terms of test error, demonstrating that the incorporation of convolutional processing greatly enhances both model accuracy and generalization.
Figure 7: (a) Comparison between partial outputs ($H_{y}$, from $t=5\Delta t$ to $22\Delta t$) of the DA-FNO model and FDTD for a single sample. (b) The last two time-step results in (a) with individually tailored colormaps. (c) Comparative scatter plot of the normalized energy for all test samples of the DA-FNO model and FDTD. The black dashed line represents $y=x$.
Based on the above experiment, the DA-FNO model with $\theta=90\%$ is employed and trained on the entire dataset, with $\delta=10^{-4}$ as the convergence criterion. The final results show a average relative $L_{1}$ error of 8.43% on the training set and 9.82% on the testing set. A partial output (from $t=5\Delta t$ to $t=22\Delta t$) of one sample with a relative $L_{1}$ error of 9.86% is illustrated in Fig. 7(a). A TM-polarized plane wave interacts with the geometry upon incidence. As the wave propagates out of the simulation domain, the energy within the region gradually decays. It can be observed from the figure that the results obtained by FDTD and the DA-FNO model are highly similar. For clearer visual comparison, the last two time-step results shown in Fig. 7(a) are re-visualized in Fig. 7(b) using individually tailored colormaps. Despite the presence of numerous high-frequency variations in the figure, the outputs of FDTD and the model exhibit strong similarity. Moreover, we compare the normalized energy at each time step of the outputs from the model and FDTD across all test samples, as shown in the Fig. 7(c). Each point represents a specific time step of a test sample. Comparison with the reference line $y=x$ reveals that the model-predicted energies closely match those computed by the FDTD in the low-energy regime. In the intermediate range, the model tends to overestimate the energy, while in the high-energy regime, it slightly underestimates it, indicating a larger variation in the model’s predictions. Despite the variation, the predicted values remain well-aligned with the reference, demonstrating an overall acceptable level of consistency between the two approaches. This consistency ensures that the model’s predictions are sufficiently reliable to allow the simulation to terminate at the correct time steps based on energy monitoring.
Figure 8: Variation of the Ez value at the central point of the simulation domain predicted by the DA-FNO model and FDTD under different geometries. The blue lines represent the FDTD predictions, and the orange dots represent the DA-FNO model predictions.
In addition, Fig. 8 presents the variations of the $E_{z}$ field at the central point of the simulation domain for 12 random samples, with the geometry of each sample displayed in the bottom-right corner of each subplot. Across various randomized geometries and simulation durations, the predictions of the DA-FNO model (orange scatter points) align closely with that of the FDTD (blue curves), indicating high accuracy in capturing temporal waveforms. Similarly, since the temporal waveform at each point is available, it is also feasible to derive corresponding frequency-domain information or energy distributions. The accuracy of such calculations depends on the temporal resolution $\Delta t_{model}$ of the simulation. A coarse time step may lead to inaccuracies in the computed frequency-domain representations, while a finer time step yields more precise results. Overall, given access to time-domain data, it is possible to obtain frequency-domain characteristics through appropriate post-processing techniques.
[…]
Figure 9: Average relative $L_{1}$ error (%) at different central wavelengths and convergence criteria.
Table 1: Average Relative $L_{1}$ Error (%) versus Central Wavelengths ($\lambda_{0}/\text{nm}$) and Convergence Criteria ($\delta$)
$\delta\backslash\lambda_{0}$ 1530 1535 1540 1545 1550 1555 1560 1565 1570
$10^{-2}$ 7.77 7.64 7.55 7.50 7.50 7.54 7.63 7.77 7.84
$10^{-3}$ 8.65 8.51 8.41 8.34 8.33 8.36 8.44 8.57 8.73
$10^{-4}$ 10.17 10.02 9.91 9.84 9.82 9.83 9.90 10.01 10.16
The DA-FNO model is capable of overcoming the CFL constraint, enabling faster simulations. To evaluate simulation speed, we test the DA-FNO model with $\theta=90\%$, the in-house FDTD, and the Meep-based FDTD [25] with different numbers of samples. All experiments are conducted on an Intel(R) Xeon(R) Gold 6346 CPU (16 cores, 32 threads) and an NVIDIA GeForce RTX 4090 GPU. For conventional FDTD methods, GPU-based parallel simulation of individual samples is inefficient when the computational domain is small, as data transfer overhead between the CPU and GPU can dominate the overall runtime. Therefore, in these experiments, the FDTD approaches are implemented with CPU-based multithreading instead, with the optimal number of threads determined empirically. In contrast, the neural network-based model naturally supports batch processing due to its architectural design, which makes it well-suited for execution on GPUs. We set the batch size to 200 in the model simulation and use O2 optimization for the C++-based in-house FDTD simulation. The Meep-based FDTD simulations are executed with 16 threads to enable parallel processing.
As shown in Table 2, the simulation speed of the DA-FNO model increases with the number of samples when the sample size is below the batch size of 200, and stabilizes at approximately 0.034 s/sample when the sample size exceeds 200. This indicates that the DA-FNO model exhibits greater computational efficiency in large-sample scenarios. In contrast, the simulation speeds of the in-house FDTD and Meep FDTD remain nearly constant at approximately 0.3 s/sample and 0.254 s/sample, respectively, regardless of the sample size. Overall, when the sample size exceeds 200, the DA-FNO model achieves a speedup of approximately $7.5\times$ compared to the multithreaded Meep FDTD simulation.
Table 2: Simulation Time (s): DA-FNO vs. FDTD across Sample Sizes
Sample Size DA-FNO in-house FDTD Meep FDTD
50 2.20 14.94 12.78
100 3.87 31.70 25.60
200 6.81 61.37 50.56
300 10.28 96.19 75.77
4 Conclusion
In this work, we propose a novel Fourier neural operator tailored for photonic simulation, termed the Data-Aware Fourier Neural Operator (DA-FNO). The DA-FNO modifies the vanilla Fourier layer by incorporating convolutional operations to capture the interactions among different components of the electromagnetic field. Furthermore, it introduces a data-aware mode selection strategy within the Fourier layer, enabling adaptive selection of Fourier modes based on data characteristics, which enhances the capability for long-term electromagnetic simulations. By embedding the DA-FNO within an auto-regressive framework, we develop a novel neural-operator-based model capable of simultaneously predicting the temporal evolution of electromagnetic fields and generalizing across complex and randomized geometries. Moreover, it demonstrates strong wavelength generalization across the optical communication C-band. The DA-FNO model iteratively outputs all field components while monitoring energy variations to automatically terminate the simulation upon convergence. It achieves a speedup of $7.5\times$ compared to the conventional FDTD method, with an accuracy of 92.5%. Given its ability to predict the temporal evolution of electromagnetic fields, generalize across randomized geometries and the optical C-band, future work will focus on applying this FDTD surrogate model to the inverse design of photonic devices with time-dependent characteristics, such as those exhibiting dispersion. The model’s computational efficiency is particularly advantageous in the early stages of inverse design, where a large number of candidate configurations must be explored and high precision is not yet critical. We believe that the DA-FNO model offers a novel approach for photonic simulations and holds great potential as a powerful tool for the inverse design of devices with time-dependent characteristics.
Acknowledgment
This work is supported by High-performance Computing Platform of Peking University.